# SePPO: Semi-Policy Preference Optimization for Diffusion Alignment

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) methods are emerging as a way to fine-tune diffusion models (DMs) for visual generation. However, commonly used on-policy strategies are limited by the generalization capability of the reward model, while off-policy approaches require large amounts of difficult-to-obtain paired human-annotated data, particularly in visual generation tasks. To address the limitations of both on- and off-policy RLHF, we propose a preference optimization method that aligns DMs with preferences without relying on reward models or paired human-annotated data. Specifically, we introduce a Semi-Policy Preference Optimization (SePPO) method. SePPO leverages previous checkpoints as reference models while using them to generate on-policy reference samples, which replace “losing images” in preference pairs. This approach allows us to optimize using only off-policy “winning images.” Furthermore, we design a strategy for reference model selection that expands the exploration in the policy space. Notably, we do not simply treat reference samples as negative examples for learning. Instead, we design an anchor-based criterion to assess whether the reference samples are likely to be winning or losing images, allowing the model to selectively learn from the generated reference samples. This approach mitigates performance degradation caused by the uncertainty in reference sample quality. We validate SePPO across both text-to-image and text-to-video benchmarks. SePPO surpasses all previous approaches on the text-to-image benchmarks and also demonstrates outstanding performance on the text-to-video benchmarks. Code will be released in https://github.com/DwanZhang-AI/SePPO.
@@ 1 Introduction
Text-to-visual models have become a crucial component of the AIGC (AI-generated content) industry, with the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020; Kingma et al., 2021) being the most widely used technology. However, current pre-trained text-to-visual models often fail to adequately align with human requirements. As a result, recent works (Liang et al., 2024a; Black et al., 2024; Wallace et al., 2024; Yang et al., 2024) adopt reinforcement learning (RL)-based approaches as the post-training process to better satisfy human needs, namely reinforcement learning from human feedback (RLHF).
These methods can generally be divided into two categories: on-policy methods and off-policy methods. Similar to large language models (LLMs), in the post-training of diffusion models, on-policy models use a reward model (RM) to score output and then backpropagate the policy gradients based on the scoring results. A typical approach is DDPO (Black et al., 2024), which utilizes vision language models as the reward model to improve the prompt-image alignment. Even though on-policy method is approved to be helpful in the Natural Language Processing (NLP) domain (Dong et al., 2024). However, aside from the existing issues with on-policy methods, such as reward hacking, which can lead to model collapse (Denison et al., 2024), in the text-to-visual task, despite the availability of numerous evaluation models, it is difficult to find a solution that can provide comprehensive feedback on all aspects of the visual content (Kim et al., 2024). Additionally, constructing an effective and efficient reward model is extremely challenging and heavily dependent on the collection of costly paired feedback data.
Another approach is to utilize a fixed set of preference data (generated by human or other models) as training data, which is called off-policy method. This method allows the trained model to achieve a distribution similar to that of the preference data. An example is Diffusion-DPO (Wallace et al., 2024), which uses the Pick-a-Pic (Kirstain et al., 2023) dataset that contains paired preference image data generated by various models with human ratings. Apparently, off-policy methods also depend on human feedback data with positive and negative sample pairs, which requires additional efforts, and their results are usually inferior to those of on-policy methods (Tang et al., 2024). Thus, in this paper, our aim is to build a preference optimization method that can mitigate the issues of on-policy and off-policy, allowing the diffusion model to align with preferences without using a reward model or paired human-annotated data.
To enable models to learn preferences from human feedback, the construction of positive and negative examples is crucial. However, in most datasets, there is usually only one sample in each data point, which typically serves as a positive example. Therefore, in the absence of a reward model, our initial consideration is how to construct appropriate negative samples. Several previous works, such as SPIN-Diffusion (Yuan et al., 2024a) and DITTO (Shaikh et al., 2024), utilize the previous checkpoints to generate the so-called “losing” samples and then use preference optimization for model training. However, these methods cannot guarantee that the samples generated from the previous checkpoints are necessarily “losing” samples relative to the current model.
To address this issue, we propose a method called Semi-Policy Preference Optimization (SePPO). In our method, the positive samples are sampled from the supervised fine-tuning (SFT) dataset. In order to obtain sufficiently good negative samples that are not too far from the current model distribution, the “losing” samples are generated by the reference model. Unlike SPIN-Diffusion, where the reference model is set to use the latest checkpoint, and DITTO, where the reference model is the initial model. In our SePPO, the reference models are sampled from all the previous checkpoints. To be specific, we first study the selection strategies for the reference model, conducting experiments using three different approaches: (1) always selecting the initial checkpoint as the reference model, (2) selecting the checkpoint saved from the previous iteration, and (3) randomly selecting from all previous checkpoints. We found that as the number of training steps increases, compared to always using the initial checkpoint, selecting the checkpoint from the previous iteration makes the reference model less prone to overfitting and yields better results. Moreover, randomly selecting from all previous checkpoints produces similar results to selecting the checkpoint from the previous iteration and leads to a more stable training process overall.
In addition, to determine whether the samples generated from the reference model are genuinely “losing” images or “winning” images relative to the current model. We further design a strategy to determine whether the sampled examples are positive or negative samples, namely Anchor-based Adaptive Flipper (AAF). If we have a winning data point for the model to learn from and the reference model has a higher probability than the current model to generate this winning data point, then the probability of sampling a winning data point from the reference model distribution will be greater than that of generating a losing data point. In other words, in this case, the “losing images” are not truly negative samples for the current model. This could negatively affect the model if we continue to use the direct preference optimization (DPO) (Rafailov et al., 2024) or SPIN (Yuan et al., 2024a) loss functions. Therefore, we design a strategy where, if the reference model has a higher probability than the current model of generating the winning data point, the model will learn from both the winning data point and the samples generated by the reference model. This not only helps avoid the negative effects of incorrectly judging sample quality but also increases the chances of the model outperforming the results of SFT to some extent.
In summary, the main contributions of this work are as follows:
@@
    We design an iterative preference optimization method called Semi-Policy Preference Optimization (SePPO). Our model can achieve preference alignment without human annotation and a reward model, which reduces labor cost and infrastructure burden;
@@
    We develop a strategy that first samples the reference model in each iteration, which enables to expand the space of policy exploration. We then design a criterion to evaluate the quality of generated responses and adjust the preference optimization based on this quality information, a process termed Anchor-based Adaptive Flipper (AAF);
@@
    SePPO exceeds all previous optimization methods on the text-to-image benchmark, and its effectiveness has also been validated on the text-to-video datasets.
@@ 2 Related Work
Self-Improvement. Self-improvement methods use iterative sampling to improve pre-trained models, which has the potential to eliminate the need for an expert annotator such as a human or more advanced models. A recent work INPO (Zhang et al., 2024) directly approximates the Nash policy (the optimal solution) of a two-player game on a human-annotated preference dataset. Concurrently, the demonstrated self-improvement methods (Shaikh et al., 2024; Chen et al., 2024) align pre-trained models only using supervised learning dataset (one demonstrated response for each prompt) without using reward models and human-annotated data pairs. Specifically, it takes the demonstrated response as the winner and generates the loser by reference models. DITTO (Shaikh et al., 2024) fixes the reference model as the initial model and works well on small datasets. SPIN (Chen et al., 2024) takes the latest checkpoint as the reference model and uses it to generate responses in each iteration. However, these approaches have the following shortcomings. First, they only focus on transformer-based models for the text modality. Second, the selection of reference models is fixed, which has limited space for policy exploration. Third, either a human-annotated preference dataset is necessarily required or they are built on a strong assumption that the demonstrated responses are always preferred to the generated responses. In fact, the responses generated by the models are not necessarily bad and do not always need to be rejected.
RM-Free PO for Diffusion Model Alignment. DPO-style methods (Rafailov et al., 2024) use a closed-form expression between RMs and the optimal policy to align pre-trained models with human preference. Thus, no RM is needed during the training process. Diffusion-DPO is first proposed in Wallace et al. (2024). Based on offline datasets with labeled human preference, it shows promising results on text-to-image diffusion generation tasks. Diffusion-RPO (Gu et al., 2024) considers semantic relationships between responses for contrastive weighting. MaPO (Hong et al., 2024) uses margin-aware information, and removes the reference model. However, only off-policy data is considered and human-annotated data pairs are necessarily required with high labor costs.
@@ 3.1 Denoising Diffusion Probabilistic Model
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})~{}={}$ $\mathop{\prod}_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),~{}~{}~{}~{}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t}|\sqrt{\alpha}\mathbf{x}_{t-1},(1-\alpha)\mathbf{I}),$ (1)
@@
where $t\sim\mathcal{U}(1,T)$ is uniformly distributed on the integer interval $[1,T]$, $w_{t}=\frac{(1-\alpha)^{2}\alpha^{t-1}}{2\sigma_{t}^{2}(1-\alpha^{t})^{2}}$ is a weighting scalar, and $\sigma_{t}^{2}=\frac{(1-\alpha)(1-\alpha^{\frac{t-1}{2}})}{1-\alpha^{t}}$ is the variance of the additive noise. $T$ as a prespecified constant in Kingma et al. (2021) is ignored in the loss function, because it has no contribution to training. In practice, $w_{t}$ is usually set to a constant (Song & Ermon, 2019).
@@ 3.2 Direct Preference Optimization
Given a prompt or condition $\mathbf{c}$, the human annotates the preference between two results as $\mathbf{x}^{w}\succ\mathbf{x}^{l}$. After preference data collection, in conventional RLHF, we train a reward function based on the Bradley-Terry (BT) model. The goal is to maximize cumulative rewards with a Kullback–Leibler (KL) constraint between the current model $\pi_{\theta}$ and a reference model $\pi_{\rm ref}$ (usually the initial model) as follows
@@
where $r(\cdot)$ is the reward function induced from the BT model.
@@
$\mathcal{L}_{\rm DPO}(\theta)~{}={}$ $\mathop{\mathbb{E}}_{(\mathbf{x}^{w},\mathbf{x}^{l},\mathbf{c})\sim\mathcal{D}}\Big{[}-\log{\sigma}\Big{(}\beta\log\frac{\pi_{\theta}(\mathbf{x}^{w}|\mathbf{c})}{\pi_{\rm ref}(\mathbf{x}^{w}|\mathbf{c})}-\beta\log\frac{\pi_{\theta}(\mathbf{x}^{l}|\mathbf{c})}{\pi_{\rm ref}(\mathbf{x}^{l}|\mathbf{c})}\Big{)}\Big{]},$ (4)
@@
(a) ${\rm ref}=0$
(b) ${\rm ref}=k-1$
(c) ${\rm ref}=[0,k-1]$ (ours)
Figure 1: Illustration of the exploration area. In the $k$-th ($k=3$) iteration, the green area satisfies the current KL constraint, while the blank area does not.
4 Iterative Alignment for Diffusion Models
4.1 Iterative Alignment from Semi-policy Preference Optimization
When we have supervised fine-tuning data ($\mathbf{c},\mathbf{x}^{w}$), we consider using previous checkpoints to construct preference pairs. Specifically, in our method, we use the reference model for the sampling distribution to ensure high-quality generation of reference samples $\mathbf{x}^{\rm ref}$. However, how to select an appropriate reference model remains a challenge.
In RL, equation 3 is a surrogate loss (Heess et al., 2017; Vieillard et al., 2020) of the hard KL-constrained target
$\max_{\pi_{\theta}}$ $\mathop{\mathbb{E}}_{\mathbf{c}\sim\mathcal{D},\atop\mathbf{x}\sim\pi_{\theta}(\cdot|\mathbf{c})}\Big{[}r(\mathbf{x},\mathbf{c})\Big{]},$ (5)
${\rm s.t.}{}$ $\mathop{\mathbb{E}}_{\mathbf{c}\sim\mathcal{D},\atop\mathbf{x}\sim\pi_{\theta}(\cdot|\mathbf{c})}\Big{[}D_{\rm KL}\big{(}\pi_{\theta}(\mathbf{x}|\mathbf{c})||\pi_{\rm ref}(\mathbf{x}|\mathbf{c})\big{)}\Big{]}\leq\delta,$
where the training path in the policy space (a family of parametric functions) can be visualized in a ball with radius $\delta$ as shown in Figure 1. $\delta$ as a hard constraint radius is roughly proportional to $1/\beta$. Assume that we save $K$ checkpoints in total during the training process. In the conventional DPO setting, the reference model is fixed to the initial checkpoint (${\rm ref}=0$). However, this setting may have a limited and fixed exploration area controlled by the hyperparameter $\beta$. To better study the problem, in the $k$-th iteration, we design and test three sampling strategies for the reference model as follows:
  • The reference model is sampled from the initial model, and is denoted as ${\rm ref}=0$.
  • The reference model is sampled from the checkpoint saved in the last iteration, and is denoted as ${\rm ref}=k-1$.
  • The reference model is randomly sampled from all previously saved checkpoints, and is denoted as ${\rm ref}=[0,k-1]$.
The training algorithm is given in Algorithm 1. We omit $\mathbf{c}$ for the sake of concision. In each iteration, the loss function adapted from diffusion DPO is as follows
$\mathcal{L}(\theta;\mathbf{x}_{0}^{w},\mathbf{x}_{0}^{\rm ref})~{}=~{}\mathop{\mathbb{E}}_{\epsilon^{w},\epsilon^{\rm ref},t}\Big{[}-\log{\sigma}\Big{(}-\beta Tw_{t}($ $\|\epsilon_{\theta}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}-\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{w})-\mathbf{\epsilon}^{w}\|^{2}$ (6)
$-\|\epsilon_{\theta}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2}+\|\epsilon_{\rm ref}(\mathbf{x}_{t}^{\rm ref})-\mathbf{\epsilon}^{\rm ref}\|^{2})\Big{)}\Big{]},$
where $t\sim\mathcal{U}(1,T)$, $\epsilon^{w},\epsilon^{\rm ref}\stackrel{{\scriptstyle\rm iid}}{{\sim}}\mathcal{N}(0,\mathbf{I})$, and $\mathbf{x}_{t}^{w,{\rm ref}}=\sqrt{\alpha^{t}}\mathbf{x}_{0}^{w,{\rm ref}}+\sqrt{1-\alpha^{t}}\epsilon^{w,{\rm ref}}$. Notably, in equation 6, the reference image $\mathbf{x}_{0}^{\rm ref}$ is generated from the reference model ${\theta}_{\rm ref}$, which is on-policy learning.
To validate the effectiveness of ${\rm ref}=[0,k-1]$, we use the Pick-a-Pic dataset and the stable diffusion 1.5 (SD-1.5) model to conduct all experiments for the ablation study. In experiments, we save a checkpoint every $30$ updates ($K=7$ iterations). The experimental results are shown in Figure 2. In the beginning, the performances of all models are improved. However, as the number of steps $T$ increases, the performance of the model with ${\rm ref}=0$ starts to decrease and quickly exhibits overfitting. The model with ${\rm ref}=k-1$ shows unstable performances. Due to the fact that when all reference samples are generated from the last checkpoint in the current round of training, the model is prone to fall into non-global optima, leading to instability in the training process Florensa et al. (2018). Therefore, considering both model performance and stability, our method uses ${\rm ref}=[0,k-1]$ to generate reference samples.
It is noteworthy that our positive samples $\mathbf{x}_{0}^{w}$ are collected off-policy from demonstrations, while the reference samples $\mathbf{x}_{0}^{\rm ref}$ are collected on-policy from the saved policies. Therefore, we term our approach Semi-Policy Preference Optimization (SePPO).

### v2

# Bridging SFT and DPO for Diffusion Model Alignment with Self-Sampling Preference Optimization

@@ Abstract
Existing post-training techniques are broadly categorized into supervised fine-tuning (SFT) and reinforcement learning (RL) methods; the former is stable during training but suffers from limited generalization, while the latter, despite its stronger generalization capability, relies on additional preference data or reward models and carries the risk of reward exploitation. In order to preserve the advantages of both SFT and RL – namely, eliminating the need for paired data and reward models while retaining the training stability of SFT and the generalization ability of RL – a new alignment method, Self-Sampling Preference Optimization (SSPO), is proposed in this paper. SSPO introduces a Random Checkpoint Replay (RCR) strategy that utilizes historical checkpoints to construct paired data, thereby effectively mitigating overfitting. Simultaneously, a Self-Sampling Regularization (SSR) strategy is employed to dynamically evaluate the quality of generated samples; when the generated samples are more likely to be winning samples, the approach automatically switches from DPO (Direct Preference Optimization) to SFT, ensuring that the training process accurately reflects the quality of the samples. Experimental results demonstrate that SSPO not only outperforms existing methods on text-to-image benchmarks, but its effectiveness has also been validated in text-to-video tasks. We validate SSPO across both text-to-image and text-to-video benchmarks. SSPO surpasses all previous approaches on the text-to-image benchmarks and demonstrates outstanding performance on the text-to-video benchmarks.
@@ 1 Introduction
Text-to-visual models (Huang et al., 2025; Li et al., 2024c) have become a crucial component of the AIGC (AI-generated content) industry, with the denoising diffusion probabilistic model (DDPM) (Ho et al., 2020; Kingma et al., 2021) being the most widely used technology. However, current pre-trained text-to-visual models often fail to adequately align with human requirements. As a result, post-training techniques (Liang et al., 2024a; Black et al., 2024; Wallace et al., 2024; Yang et al., 2024; Li et al., 2024a; b) are widely used to align pre-trained models to better satisfy human needs.
Figure 1: Illustration of SFT, DPO, and the proposed SSPO. $x^{w}$ is the winning samples, and $x^{l}$ is the losing samples. $x^{\rm rand\_w}$ and $x^{\rm rand\_l}$ are the samples generated by randomly sampled checkpoints. SSPO can switch between SFT and DPO based on the current state of the datapoint.
Currently, post-training techniques that are widely used can generally be categorized into two main types: supervised fine-tuning (SFT) and reinforcement learning (RL) methods. However, each approach has its advantages and limitations. SFT requires only target distribution data for training and exhibits greater stability in the training process compared to RL. However, its generalization ability is significantly inferior to that of RL methods (Chu et al., 2025). Conversely, common RL methods, despite their strong generalization capabilities, often necessitate additional data or model constraints. For instance, Diffusion-DPO (Wallace et al., 2024) employs a fixed set of preference data (generated by humans or other models) as training data. Methods such as DDPO (Black et al., 2024), CDPO (Croitoru et al., 2024), and SPO (Liang et al., 2024b) utilize a reward model (RM) (Escontrela et al., 2024) to score outputs and subsequently backpropagate policy gradients based on the scoring results. This not only incurs additional computational costs but also introduces the risk of reward hacking (Denison et al., 2024). Moreover, in the context of text-to-visual tasks, despite the existence of numerous evaluation models, finding a solution capable of providing comprehensive feedback on all aspects of visual content remains challenging (Kim et al., 2024). Also, constructing an effective and efficient reward model is particularly difficult, as it heavily relies on the collection of costly paired feedback data. This naturally raises the question of whether it is possible to develop an approach that combines the advantages of both SFT and RL – eliminating the need for paired preference data or reward models required by RL methods, while retaining the training stability of SFT and achieving a level of generalization comparable to RL. In this work, as shown in Figure. 1, we introduce a method called Self-Sampling Preference Optimization (SSPO) to tackle this challenge.
Since only SFT data (“winning” samples) can be used, we need to construct appropriate new samples to complete the DPO training process. Some previous methods have utilized sampling from historical distributions to construct new samples, such as Experience Replay (ER) Lin (1992); Schmitt et al. (2018); Pinto et al. (2017). ER improves learning efficiency and effectiveness by storing and reusing past experiences (historical data stored in the buffer). This concept is also reflected in SPIN-Diffusion Yuan et al. (2024a), where samples are drawn from the latest checkpoint to generate negative samples. However, such methods face two major challenges:
  • What kind of historical distribution is worth sampling?
  • Are the sampled data truly negative samples?
To address the first issue, we employ historical checkpoints as the minimal sampling unit of the Experience Replay Distribution (ERD). Specifically, we conduct experiments using three different approaches to identify the optimal sampling strategy for ERD: (1) consistently selecting the initial checkpoint as the ERD, (2) selecting the last checkpoint saved from the previous iteration, and (3) uniform sampling from all previous checkpoints. Our findings indicate that as the number of training steps increases, compared to always utilizing the last checkpoint, selecting the checkpoint from the previous iteration can mitigate the risk of overfitting and yields superior performance. Furthermore, randomly selecting from all previous checkpoints produces comparable results to selecting the checkpoint from the previous iteration while contributing to a more stable overall training process. Therefore, we refer to this final strategy as Random Checkpoint Replay (RCR), which enables the model to learn from historical information better, ensuring both training stability and outstanding performance.
Meanwhile, regarding the second issue, previous strategies lack any guarantee that the sampled data points are necessarily negative samples. However, in DPO, a well-defined “winning” or “losing” paired dataset is essential. Therefore, to determine whether the data generated by the reference model are genuinely “losing” or “winning” samples relative to the current policy, we propose a novel strategy called Self-Sampling Regularization (SSR). This strategy enables the classification of sampled examples as either positive or negative samples. Specifically, if a winning data point is available for the model to learn from and the reference model exhibits a higher probability than the current model of generating this winning data point, then the probability of sampling a winning data point from the reference model’s distribution will exceed that of generating a losing data point. In other words, under such conditions, the so-called “losing” samples do not constitute truly negative samples for the current model. In such cases, SSPO automatically shifts the learning paradigm from DPO to SFT, thereby preventing the model from learning biases from samples that are not genuinely “losing” when using the DPO algorithm.
In summary, the main contributions of this work are:
@@
    We design a new preference algorithm SSPO that, for the first time, integrates the advantages of SFT and DPO, achieving superior optimization results.
@@
    We design the Random Checkpoint Replay strategy to construct paired data, then experimentally and theoretically analyze it as an effective solution in sampling of the experience replay distribution.
@@
    We further design a criterion termed Self-sampling Regularization to evaluate the quality of the generated responses, which subsequently determines whether the current sample is used for DPO or SFT.
  4. 4.
    SSPO outperforms all previous optimization methods on the text-to-image benchmarks, and is also effective on the text-to-video benchmarks.
@@ 2 Related Work
Self-Improvement. Self-improvement methods use iterative sampling to improve pre-trained models, which has the potential to eliminate the need for an expert annotator such as a human or more advanced models. A recent work INPO (Zhang et al., 2025) directly approximates the Nash policy (the optimal solution) of a two-player game on a human-annotated preference dataset. Concurrently, the demonstrated self-improvement methods (Shaikh et al., 2024; Chen et al., 2024) align pre-trained models only using supervised learning dataset (one demonstrated response for each prompt) without using reward models and human-annotated data pairs. Specifically, it takes the demonstrated response as the winner and generates the loser by reference models. DITTO (Shaikh et al., 2024) fixes the reference model as the initial model and works well on small datasets. SPIN (Chen et al., 2024) takes the latest checkpoint as the reference model and uses it to generate responses in each iteration. However, these approaches have the following shortcomings. First, they only focus on transformer-based models for the text modality. Second, the selection of reference models is fixed, which has limited space for policy exploration. Third, either a human-annotated preference dataset is necessarily required or they are built on a strong assumption that the demonstrated responses are always preferred to the generated responses. In fact, the responses generated by the models are not necessarily bad and do not always need to be rejected.
RM-Free PO for Diffusion Model Alignment. DPO-style methods (Rafailov et al., 2024) use a closed-form expression between RMs and the optimal policy to align pre-trained models with human preference. Thus, no RM is needed during the training process. Diffusion-DPO is first proposed in Wallace et al. (2024). Based on offline datasets with labeled human preference, it shows promising results on text-to-image diffusion generation tasks. Diffusion-RPO (Gu et al., 2024) considers semantic relationships between responses for contrastive weighting. MaPO (Hong et al., 2024) uses margin-aware information, and removes the reference model. Diffusion-NPO (Wang et al., 2025) further adds label smoothing on Diffusion-DPO. However, only off-policy data is considered and human-annotated data pairs are necessarily required with high labor costs.
@@ 3.1 Denoising Diffusion Probabilistic Model
$q(\mathbf{x}_{1:T}|\mathbf{x}_{0})~{}={}$ $\mathop{\prod}_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1}),\ \ q(\mathbf{x}_{t}|\mathbf{x}_{t-1})~{}=~{}\mathcal{N}(\mathbf{x}_{t}|\sqrt{\alpha}\mathbf{x}_{t-1},(1-\alpha)\mathbf{I}),$ (1)
@@
where $t\sim\mathcal{U}(1,T)$ is uniformly distributed on the integer interval $[1,T]$, $w_{t}=\frac{(1-\alpha)^{2}\alpha^{t-1}}{2\sigma_{t}^{2}(1-\alpha^{t})^{2}}$ is a weighting scalar, and $\sigma_{t}^{2}=\frac{(1-\alpha)(1-\alpha^{\frac{t-1}{2}})}{1-\alpha^{t}}$ is the variance of the additive noise. $T$ as a pre-specified constant in Kingma et al. (2021) is ignored in the loss function, because it has no contribution to training. In practice, $w_{t}$ is usually set to a constant (Song & Ermon, 2019).
@@ 3.2 Direct Preference Optimization
Given a prompt or condition $\mathbf{c}$, the human annotates the preference between two results as $\mathbf{x}^{w}\succ\mathbf{x}^{l}$. After preference data collection, in conventional RLHF (Reinforcement Learning from Human Feedback), we train a reward function based on the Bradley-Terry (BT) model. The goal is to maximize cumulative rewards with a Kullback–Leibler (KL) constraint between the current model $\pi_{\theta}$ and a reference model $\pi_{\rm ref}$ (usually the initial model) as follows
@@
where $r(\cdot)$ is the reward function induced from the BT model, and $\beta$ is a hyperparameter to control the weight of the KL constraint.
@@
$\mathcal{L}_{\rm DPO}(\theta)$ $~{}=~{}\mathop{\mathbb{E}}_{(\mathbf{x}^{w},\mathbf{x}^{l},\mathbf{c})\sim\mathcal{D}}\Big{[}-\log\ \ {\sigma}\Big{(}\beta\log\frac{\pi_{\theta}(\mathbf{x}^{w}|\mathbf{c})}{\pi_{\rm ref}(\mathbf{x}^{w}|\mathbf{c})}-\beta\log\frac{\pi_{\theta}(\mathbf{x}^{l}|\mathbf{c})}{\pi_{\rm ref}(\mathbf{x}^{l}|\mathbf{c})}\Big{)}\Big{]},$ (4)
@@
4 Self-Sampling Preference Optimization
4.1 Random Checkpoint Replay
When we only have supervised fine-tuning data ($\mathbf{c},\mathbf{x}^{w}$), curating the paired data $\mathbf{x}^{\rm rand}$ becomes crucial. However, how to select an appropriate ERD remains a challenge. Moreover, unlike DPO, SSPO combines both SFT and DPO, which makes the reference model more influenced by the ERD rather than the initial model. Consequently, we adopt the ERD as our reference model. To further investigate the impact of ERD sampling, we design and evaluate three sampling strategies for the ERD as follows:
  • The ERD is sampled from the initial model, and is denoted as ${\rm ERD}=0$.
  • The ERD is sampled from the checkpoint saved in the last iteration, and is denoted as ${\rm ERD}=k-1$.
  • The ERD is uniformly sampled from all previously saved checkpoints, and is denoted as ${\rm ERD}=[0,k-1]$.